OpenStudy (chupacabraj):

What fomula(s), relationships, concept(s), should I know before attempting to solve this problem?

OpenStudy (chupacabraj):

I'm going over a sample final exam for my physics 2 class and I came across this problem but I have no idea how to go about it.

OpenStudy (mrs.ambrose614):

Can you tell me the problem I can't open the link???

OpenStudy (irishboy123):



OpenStudy (chupacabraj):

so I found a formula V = Ed

OpenStudy (irishboy123):

if you're unsure about this, look at the units. i have no idea what level you are working at but the science is \(E = - \dfrac{\Delta V}{\Delta x}\), if that makes sense. Ah yes, your post ..... that's the ticket!

OpenStudy (chupacabraj):

Kind of solve for ddelta v then

OpenStudy (irishboy123):

Reconciling our formulations.... \(\Delta V = - E \Delta x\) the minus sign is important but it follows from common sense ***key idea*** that is hidden away here is that Potential is a measure of Energy per unit charge. So force on a charge in a E field is \(F = E q\) Work done against field by moving the charge upstream is \(W = - F \cdot x = - Eqx = U\), where U is energy \(V = \dfrac{U}{q} = - Ex\) fiddly but this is IMHO the best way to get your head around this and to be able to generate useful formulae rather than having to try to remember them. the minus sign isn't too crucial either, as long as you remember that the direction of an E field is the direction in which a proton is compelled to move